# The BPgWSP test: a Bayesian Weibull Shape Parameter signal detection test for adverse drug reactions

### v1

@@ 4.3 Alopecia
Figure 10: Graphics for exploratory analysis of the occurrence of alopecia over three years. The graphic derived is based on Sauzet et al., (2013, Fig. 3).
@@
Over three years there have been 144 reports of alopecia in the group of 19812 patients (0.73%) exposed to bisphosphonates. The mean event-time is on day 424 (285). According to the histogram (Figure 10 1.a) events tend to take place slightly more often between the fifth and 13th month which roughly corresponds to the period around the first or second quarter of the observation period. Regarding the estimated hazard function the risk rises and falls within the first two years and stabilizes after that.
There are 180 alopecia cases among 39634 control patients ($0.45\%$). The event-times are slightly more dense in the first half of the observation window. The estimated hazard (Figure 10 2.b) has an almost flat trend.
@@
prior belief about prior means test outcome test outcome
@@
none $(1,1,1)$ signal -
1st quarter of OP (274 days) $(200,0.63,1)$ signal -
2nd quarter of OP (548 days) $(30,3.2,10)$ signal -
3rd quarter of OP (821 days) $(700,14,4)$
Table 9: Test results under various prior beliefs for alopecia based on exposed and control group.
Table 9 shows the BPgWSP test results for alopecia. Under the prior assumptions of no association, ADR occurrence in the beginning or in the middle of the observation period the test identifies a temporal pattern for the exposed group and none for the control group. For both groups, no signal is raised under the prior belief that alopecia risk increases at the end of the observation period.
Alopecia is a rare event. In the non-Bayesian approach in Sauzet et al., (2013) an ambiguous signal was given if the data was censored after one year only whereas here, we get a consistent signal if the prior belief does not expect cases to occur toward the three-year mark. For the control group, no signal is raised indicating that the signal raised for cases is probably not due to noise.
4.4 Carpal tunnel syndrome
Figure 11: Graphics for exploratory analysis of the occurrence of carpal tunnel syndrome over three years. The graphic derived is based on Sauzet et al., (2013, Fig. 4).
Exploratory Analysis:
Within the observation period of three years, 192 among 19817 exposed patients ($0.97\%$) experienced carpal tunnel syndrome. The day of occurrence is on average $429\;(5.9)$ which is approximately one year after drug prescription. The histogram of event-times (Figure 11 a.1) reveals that events very slightly decrease over the scope of the observation period. The hazard estimate (Figure 11 b.1) indicates a slightly higher constant risk in the first 18 months and a decrease after that.
Within the control group, carpal tunnel syndrome occurred in 244 among 39389 patients ($0.62\%$). The distribution of events looks similar to the exposed group’s plot (see Figure 7 a.2). The hazard estimate shows an incline of risk in the first year, a rather constant trend in the middle and again an increase within the last 5 months. (see Figure 11 b.2).
Signal detection test:
prior belief about prior means test outcome test outcome
none $(1,1,1)$
1st quarter of OP (274 days) $(200,0.63,1)$ signal signal
2nd quarter of OP (548 days) $(30,3.2,10)$ signal signal
3rd quarter of OP (821 days) $(700,14,4)$
Table 10: Test outcomes under various prior beliefs for carpal tunnel syndrome based on exposed and control group.
The BPgWSP test results for carpal tunnel syndrome are presented in Table 10. In two out of four prior specifications, namely when the prior belief is that the event is to be expected around the first quarter or middle of the observation period, a signal is raised for both control and exposed group. In the two other cases, a signal is raised for neither group.
The ambiguous test results occurring here and in Sauzet et al., (2013) where no signal was raised for the exposed group and a signal was raised for the control group when censored after 637 days are probably returned due to the confounding effect of the age of the participants for which the carpal tunnel syndrome is a common condition.
@@ 5 Discussion
In this work we developed the BPgWSP test that raises a signal if there is an indication of non-constant hazard over time for an AE after drug intake incorporating prior knowledge about the time of occurrence. The defined test was tuned based on a simulation study which led to a one-fits-all approach. Consequently, we applied the test with the recommended tuning to real data.
The underlying assumption of the BPgWSP test is that the hazard should be constant if there is no drug-AE association. This assumption also underlies the work of Cornelius et al., (2012) and the extension by Sauzet and Cornelius, (2022) in the frequentist framework. While this heuristic assumption can be refuted due to various confounding factors, we argue that it is a simplified model is good enough for flagging possible associations because finding causal association is not the aim of the procedure.
The simulation study enabled an initial assessment of the test performance and the conclusion of a recommended BPgWSP test tuning for application in an automated one-fits-all manner to a multitude of drug-AE combinations. Arguably, the amount of combinations of simulation scenarios and prior specifications considered within the frame of the simulation study is too limited to draw general conclusions on tuning parameters and only give advise for tuning choices. For instance only one censoring time (365) was considered for data generation and only one standard deviation (10) was considered for the prior belief specification such that insights on the expected performance in case of alternative specifications cannot be drawn. This was due to the practicality of presenting the methods. In general however, we recommend to tune the BPgWSP test with an additional simulation study adjusted for the circumstances of interest. To do so, we provide the R package BWSPsignal available on https://github.com/julia-dyck/BWSPsignal such that new values to specify the data generating characteristics (see sec. 3.1), model specifications (see sec. 3.2) and tuning parameters for the HDI+ROPE test setup (see sec. 3.3) can be easily evaluated.
In the case study, we applied the BPgWSP test to real data on bisphosphonates with adverse events known to be ADRs caused by the drug (headache, musculoskeletal pain) and adverse events not identified as ADR (alopecia, carpal tunnel syndrome). The BPgWSP test tuned based on the simulation results was able to raise a signal under varying prior beliefs for the known ADRs. It returned ambiguous test results for adverse events not categorized as ADRs under some prior belief specifications. The performance of the BPgWSP test depends on whether the observation period encompasses the expected event-time and on the rareness of the adverse event. This is an additional argument for the need of more differentiated and context specific tuning of the BPgWSP test. Simulations representing the data scenarios better could improve the test performance in the case study applications.
The data used for the case study is a historical dataset derived from the THIN database (THIN,, 2024). It would be desirable to test the newly developed methods with a larger set of drug-AE pairs. However, EHR data are not easily available for method development due todata privacy regulations and costs. There are attempts to create synthetic, yet realistic datasets. An example is the SYNTHEA database (Walonoski et al.,, 2017). Even though the SYNTHEA database is based on a complex data-generating process including modules for a set of health conditions and medications that occur with probabilities and prevalences reflecting reality, the generating process does not include AEs. We think that further development of SYNTHEA and similarly structured databases needs to be supported even more extensively to facilitate the development of statistical methods for pharmacoepidemiology and other fields.
6 Conclusion
Signal detection is an important component of pharmacovigilance to monitor drugs post market release. To improve further development of statistical signal detection methods, we combine two characteristics of existing approaches in the literature, the use of longitudinal data from EHRs and the incorporation of prior knowledge about the drug under investigation, in one test.
The Bayesian WSP test is based on the constant-hazard assumption and consists of fitting a Bayesian PgW model with a prior distribution for the PgW parameters reflecting the prior belief about the drug-AE relationship, as well as a tuned HDI+ROPE test based on the posterior shape parameters. We showed how to tune the parameters specif using a simulation study, and illustrated it on data about bisphosphonate prescriptions.

### v3

@@ 4.3 Alopecia
Figure 10: Graphics for exploratory analysis of the occurrence of alopecia over three years. The graphic derived is based on Sauzet et al., Fig. 3. 32
@@
Over three years, there have been 144 reports of alopecia in the group of 19812 patients (0.73%) exposed to bisphosphonates. The mean event-time is day 424 $(sd=285)$. According to the histogram (Figure 10 1.a), events take place slightly more often between the fifth and 13th month which roughly corresponds to the period around the first or second quarter of the OP. Regarding the estimated hazard function, the risk rises and falls within the first two years and stabilizes after that.
There are 180 alopecia cases among 39634 control patients ($0.45\%$). The event-times are slightly denser in the first half of the observation window. The estimated hazard (Figure 10 2.b) has an almost flat trend.
prior belief about a priori $E[t]$ prior means test outcome test outcome
none - $(1,1,1)$ signal -
1st quarter of OP 274 days $(200,0.63,1)$ signal -
2nd quarter of OP 548 days $(30,3.2,10)$ signal -
3rd quarter of OP 821 days $(700,14,4)$ - -
Table 8: Test results under various prior beliefs for alopecia based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
@@
Under the prior assumptions of no association, ADR occurrence in the beginning or in the middle of the OP the test identifies a temporal pattern for the exposed group and none for the control group. For both groups, no signal is raised under the prior belief that alopecia risk increases at the end of the OP (see Table 8).
Alopecia is a rare event not certainly identified as an ADR of bisphosphonates at the current state of research to our knowledge. The same uncertainty is reflected in the BPgWSP test results. Under an ADR-suspecting prior, no signal is raised for the control group and a signal flagged for the exposed group indicating that the signal might not be raised due to noise. Even with larger posterior samples under the none-prior, no signal was identified. This indicates that more iterations do not necessarily lead to a higher signal rate.
4.4 Carpal tunnel syndrome
Figure 11: Graphics for exploratory analysis of the occurrence of carpal tunnel syndrome over three years. The graphic derived is based on Sauzet et al., Fig. 4. 32
Exploratory Analysis:
Within the OP of three years, 192 among 19817 exposed patients ($0.97\%$) experienced carpal tunnel syndrome. The day of occurrence is on average 429(sd=290)429\;(sd=290) which is approximately one year after drug prescription. The histogram of event-times (Figure 11 a.1) reveals that events very slightly decrease over the scope of the OP. The hazard estimate (Figure 11 b.1) indicates a slightly higher constant risk in the first 18 months and a decrease after that.
Within the control group, carpal tunnel syndrome occurred for 244 among 39389 patients ($0.62\%$). The distribution of events looks similar to the exposed group’s plot (see Figure 8 a.2). The hazard estimate shows an incline of risk in the first year, a rather constant trend in the middle and again an increase within the last 5 months. (see Figure 11 b.2).
Signal detection test:
In two out of four prior specifications, namely when the prior belief is that the event is to be expected around the first quarter or middle of the OP, a signal is raised for both the control and exposed group. In the two other cases, a signal is raised for neither group (see Table 9).
Carpal tunnel syndrome is not a recognized ADR to our knowledge. The ambiguous test results assigning signals to exposed and control group under a subset of prior specifications are probably returned due to the confounding effect of the age of the participants for which the carpal tunnel syndrome is a common condition. This suggests that sample stratification might be an important preparation before BPgWSP testing in some applications.
prior belief about a priori $E[t]$ prior means test outcome test outcome
@@ 4.4 Carpal tunnel syndrome
none - $(1,1,1)$ - -
1st quarter of OP 274 days $(200,0.63,1)$ signal signal
2nd quarter of OP 548 days $(30,3.2,10)$ signal signal
3rd quarter of OP 821 days $(700,14,4)$ - -
Table 9: Test outcomes under various prior beliefs for carpal tunnel syndrome based on exposed and control group. Prior standard deviation is set to 10 for all parameters in all prior belief cases.
@@ 5 Discussion
In this work we developed the BPgWSP test that raises a signal if there is an indication of non-constant hazard over time for an AE after drug intake incorporating prior knowledge about the time of occurrence. The defined test was tuned based on a simulation study which led to a one-fits-all approach.
An important aspect of our proposed BPgWSP test is the role of prior information in shaping its performance. The BPgWSP test performs very well under correct prior assumptions and outperforms the frequentist approach. If the prior specification is wrong, the performance of the test decreases considerably which is also reflected in the case study.
We recommend using the BPgWSP test only in case of certainty about prior information, and otherwise, use the frequentist approach. In light of these results, correct prior elicitation is of importance. It is advised to follow procedures recommended by others. See O’Hagan for an introduction.35
The simulation study enabled an initial assessment of the test performance and the conclusion of a recommended BPgWSP test tuning for application in an automated one-fits-all manner to a multitude of drug-AE combinations. In general, however, we recommend tuning the BPgWSP test with an additional simulation study. For this purpose, we developed the R package BWSPsignal.26
@@
Simulation study output as well as the R package BWSPsignal and R scripts for reproducibility are available on https://github.com/julia-dyck/BWSPsignal.
The data used in the case study section was originally provided from the THIN 30 database. Please contact the corresponding author upon reasonable request.